# Query Performance Prediction using Relevance Judgments Generated by Large Language Models

### v2

@@ Abstract.
We face two main challenges: (i) excessive computational costs of judging an entire corpus for predicting a metric considering recall, and (ii) limited performance in prompting open-source LLMs in a zero-/few-shot manner. To solve the challenges, we devise an approximation strategy to predict an IR measure considering recall and propose to fine-tune open-source LLMs using human-labeled relevance judgments. Experiments on the TREC 2019–2022 deep learning tracks show that QPP-GenRE achieves state-of-the-art QPP quality for both lexical and neural rankers.
Query performance prediction, Large language models, Relevance judgments
copyright: rightsretainedjournalyear: 2024journal: TOISjournalvolume: 1journalnumber: 1article: 1publicationmonth: 6ccs: Information systems Evaluation of retrieval results
@@ 1. Introduction
QPP, a.k.a. query difficulty prediction, has attracted the attention of the information retrieval (IR) community throughout the years (Arabzadeh et al., 2024; Ganguly et al., 2022; Carmel and Yom-Tov, 2010). QPP aims to estimate the retrieval quality of a search system for a query without using human-labeled relevance judgments (Faggioli et al., 2021b). Effective QPP benefits various tasks (Faggioli et al., 2023b), e.g., query variant selection (Di Nunzio and Faggioli, 2021; Scells et al., 2018; Thomas et al., 2017), IR system configuration selection (Deveaud et al., 2016; Tonellotto et al., 2013), and query-specific pool depth prediction (Ganguly and Yilmaz, 2023) to reduce human relevance judgment costs.
@@
QPP methods can be applied in various domains and scenarios (Faggioli et al., 2023b). We are usually concerned with the predicted retrieval quality w.r.t. various IR measures across different scenarios, e.g., our emphasis might be on precision (Faggioli et al., 2023d, 2021a) for conversational search and on recall for legal search (Tomlinson et al., 2007). However, existing QPP approaches typically predict only a single real-valued score that indicates the retrieval quality for a query (Ganguly et al., 2022) and do not require the predicted score to approximate a specific IR evaluation measure (Arabzadeh et al., 2023; Singh et al., 2023; Tao and Wu, 2014; Shtok et al., 2012; Zhou and Croft, 2007). This results in two key limitations: (i) While predicted performance scores have been shown to correlate with some IR evaluation metrics (Datta et al., 2022c; Ganguly et al., 2022), relying on a single value to represent different IR evaluation measures leads to a “one size fits all” approach, which is problematic because the literature shows that some IR metrics do not correlate well and the agreement varies across scenarios and queries (Gupta et al., 2019; Jones et al., 2015). Although some studies train regression-based QPP models to predict a specific IR evaluation measure (Khodabakhsh and Bagheri, 2023; Chen et al., 2022; Datta et al., 2022a; Arabzadeh et al., 2021b; Hashemi et al., 2019), they require training separate models to predict different measures, leading to lots of storage and running costs. (ii) A single-score prediction limits the interpretability of QPP. It is insufficient to explain QPP outputs or to analyze and fix inaccurate QPP results based solely on a single score. We argue that more in-depth and interpretable insights into QPP outputs are required.
@@
We propose a QPP framework using automatically generated relevance judgments (QPP-GenRE), in which we decompose QPP into independent subtasks of automatically predicting the relevance of each item in a ranked list for a given query. QPP-GenRE comes with various advantages: (i) it allows us to directly predict any desired IR evaluation measure at no additional cost, using automatically generated relevance judgments as pseudo-labels; and (ii) the generated relevance judgments provide an explanation beyond simply gauging how difficult or easy a query is by offering information about why the query is predicted as being difficult or easy; moreover, we can translate the “QPP errors” into easily observable “relevance judgment errors,” e.g., false positives or negatives, informing potential ways of improving QPP quality by fixing observed relevance judgment errors.
@@
QPP-GenRE can be integrated with various approaches for judging relevance. Obviously, for QPP-GenRE to excel in QPP, it is crucial to equip QPP-GenRE with an approach capable of accurately generating relevance judgments. Recent studies (Rahmani et al., 2024; Ma et al., 2024; Thomas et al., 2023; Faggioli et al., 2023a; MacAvaney and Soldaini, 2023; Gilardi et al., 2023) have shown the potential effectiveness of using LLMs to generate relevance judgments. We build on them to equip QPP-GenRE with LLMs for judging relevance. However, those studies have certain limitations: several authors have prompted commercial LLMs (e.g., ChatGPT, GPT-3.5/4, GPT-4o) to generate relevance judgments (e.g., Upadhyay et al., 2024b; Rahmani et al., 2024; Upadhyay et al., 2024a; Zhang et al., 2024; Ma et al., 2024; Thomas et al., 2023; Faggioli et al., 2023a); commercial LLMs come with limitations like non-reproducibility, non-deterministic outputs and potential data leakage between pretraining and evaluation data, impeding their value in scientific research (Pradeep et al., 2023b; Zhang et al., 2023c; Pradeep et al., 2023a). Although MacAvaney and Soldaini (2023) prompt small-scale open-source language models (e.g., Flan-T5 (Chung et al., 2022) with 3B parameters) for generating relevance judgments, they focus on a setting wherein the model is already given one relevant item for each query, which does not apply to QPP as we typically do not know any relevant item for a query in advance. In this paper, we focus on the use of open-source LLMs for generating relevance judgments in a realistic setting where we lack prior knowledge of any relevant items for a query. There are only few studies (Zhang et al., 2024; Salemi and Zamani, 2024; Upadhyay et al., 2024a; Khramtsova et al., 2024) attempting to prompt open-source LLMs in this setting.
@@
We face two challenges when using QPP-GenRE for QPP: (i) predicting IR metrics considering recall, ideally, entails identifying all relevant items in the entire corpus for a query; however, using an LLM to judge the entire corpus per query is impractical due to the significant computational overhead; (ii) our experiments reveal that directly prompting open-source LLMs in a zero-/few-shot manner yields limited effectiveness in predicting relevance, resulting in limited QPP quality; this aligns with recent findings indicating limited success in prompting open-source LLMs for specific tasks (Qin et al., 2023).
@@
We integrate QPP-GenRE with a leading open-source LLM, LLaMA (Touvron et al., 2023b, a) for generating relevance judgments; LLaMA has shown competitive performance on many natural language processing (NLP) and IR tasks (Touvron et al., 2023b, a). Experiments on datasets from the TREC 2019–2022 deep learning (TREC-DL) tracks (Craswell et al., 2022, 2021, 2020, 2019) show that QPP-GenRE achieves state-of-the-art QPP quality in estimating the retrieval quality of a lexical ranker (BM25) and a neural ranker (ANCE, Xiong et al., 2021) in terms of RR@10, a precision-oriented IR metric, and nDCG@10, an IR metric considering recall. See Sections 6.1 and 6.2.
@@
Moreover, we investigate the impact of fine-tuning and the choice of LLMs on the quality of generated relevance judgments and QPP. Besides LLaMA-7B, we consider the newly-released Llama-3-8B (AI@Meta, 2024), and Llama-3-8B-Instruct (AI@Meta, 2024), and compare their fine-tuned versions against their few-shot prompted counterparts. We find that fine-tuning markedly improves the quality of relevance judgment generation and QPP for all LLMs; the performance of fine-tuned LLMs in terms of judging relevance exceeds that of a commercial LLM (GPT-3.5) (Faggioli et al., 2023a). Also, we observe that the newly released and instruction-tuned LLMs generally deliver better performance. This implies that with a more effective LLM, QPP-GenRE has the potential to be more effective at QPP. See Section 7.2.
Additionally, to show QPP-GenRE’s compatibility with other types of relevance prediction methods, we adapt a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a), into a relevance judgment generator by applying a threshold to its re-ranking scores. Our results indicate that QPP-GenRE integrated with RepLLaMA achieves high QPP quality, at the cost of tuning a proper threshold. The high QPP quality achieved by RankLLaMA demonstrates QPP-GenRE’s compatibility with other types of relevance prediction methods. See Section 7.3.
We also analyze QPP errors based on automatically generated relevance judgments, demonstrating QPP-GenRE’s interpretability. See Section 7.4.
Finally, our computational cost analysis shows that QPP-GenRE shows lower latency than some supervised QPP baselines when predicting multiple measures because multiple measures can be derived from the same set of relevance judgments. Although QPP-GenRE shows higher latency than other QPP baselines when predicting only one metric, QPP-GenRE ’s latency is still 20 times smaller than the state-of-the-art GPT-4-based listwise re-ranker (Sun et al., 2023b). See Section 7.5.
@@
Given QPP-GenRE’s high QPP quality and interpretability, it is well-suited for some knowledge-intensive professional search scenarios, e.g., legal (Tomlinson et al., 2007) or patent search (Lupu et al., 2013), where accurate QPP is prioritized, interpretable QPP results are preferred, and users may have a higher tolerance level for latency than users in web search. Plus, QPP-GenRE can be used to analyze how well a search system performs in offline settings (Faggioli et al., 2023d), where latency is not necessarily an issue.
@@
To facilitate future research, we release our data, scripts for fine-tuning/inference, sampled demonstration examples for few-shot prompting, and fine-tuned checkpoints for LLaMA-7B, Llama-3-8B, and Llama-3-8B-Instruct at https://github.com/ChuanMeng/QPP-GenRE.
@@
  • We devise an approximation strategy to predict IR measures considering recall, avoiding the cost of traversing the entire corpus to identify all relevant items for a query.
  • We fine-tune leading open-source LLMs, LLaMA-7B, Llama-3-8B, and Llama-3-8B-Instruct for the task of automatically generating relevance judgments.
@@ 2.1. Query performance prediction
A large number of unsupervised and supervised post-retrieval QPP methods have been proposed (Carmel and Yom-Tov, 2010) for predicting the performance of lexical rankers, such as query likelihood (Lafferty and Zhai, 2001) and BM25 (Robertson et al., 2009). Unsupervised QPP methods can be classified into clarity-based (Cronen-Townsend et al., 2002), robustness-based (Zhou and Croft, 2007, 2006; Aslam and Pavlu, 2007), coherence-based (Arabzadeh et al., 2021a; Diaz, 2007), and score-based (Tao and Wu, 2014; Shtok et al., 2012; Cummins et al., 2011; Pérez-Iglesias and Araujo, 2010; Zhou and Croft, 2007). More recently, a set of supervised QPP methods have been proposed (Arabzadeh et al., 2021b; Hashemi et al., 2019; Zamani et al., 2018; Datta et al., 2022a, b; Chen et al., 2022; Khodabakhsh and Bagheri, 2023). NeuralQPP (Zamani et al., 2018) and Deep-QPP (Datta et al., 2022a) are optimized from scratch. NQA-QPP (Hashemi et al., 2019) and BERT-QPP (Arabzadeh et al., 2021b) fine-tune BERT (Devlin et al., 2019) to improve QPP effectiveness. Further, Datta et al. (2022c) propose qppBERT-PL, which considers list-wise-document information, while Chen et al. (2022) propose BERT-groupwise-QPP that considers both cross-query and cross-document information. Khodabakhsh and Bagheri (2023) propose a multi-task query performance prediction framework (M-QPPF), learning document ranking and QPP simultaneously.
@@
2.2. Zero/few-shot prompting and parameter-efficient fine-tuning (PEFT) for LLMs
While fine-tuning pre-trained language models has given rise to many state-of-the-art results (Devlin et al., 2019), fully fine-tuning LLMs for a specific task on consumer-level hardware is typically infeasible (Zhu et al., 2023) because of the large number of parameters of LLMs. As a result, there are three prevailing ways to adapt LLMs to a specific task: zero-shot prompting, few-shot prompting, a.k.a.  in-context learning (ICL(Dong et al., 2022; Brown et al., 2020), and parameter-efficient fine-tuning (PEFT(Dettmers et al., 2023; Liu et al., 2022; Hu et al., 2021).

### v3

@@ Abstract.
We face two main challenges: (i) excessive computational costs of judging an entire corpus for predicting a metric considering recall, and (ii) limited performance in prompting open-source LLMs in a zero-/few-shot manner. To solve the challenges, we devise an approximation strategy to predict an IR measure considering recall and propose to fine-tune open-source LLMs using human-labeled relevance judgments. Experiments on the TREC 2019–2022 deep learning tracks and CAsT-19–20 datasets show that QPP-GenRE achieves state-of-the-art QPP quality for both lexical and neural rankers.
Query performance prediction, Large language models, Relevance judgments, Relevance prediction, Re-ranking, Conversational search
copyright: rightsretainedjournal: TOISjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3736402ccs: Information systems Evaluation of retrieval results
@@ 1. Introduction
QPP, a.k.a. query difficulty prediction, has attracted the attention of the information retrieval (IR) community throughout the years (Arabzadeh et al., 2025, 2024b, 2024a; Ganguly et al., 2022; Carmel and Yom-Tov, 2010). QPP aims to estimate the retrieval quality of a search system for a query without using human-labeled relevance judgments (Faggioli et al., 2021b). Effective QPP benefits various downstream applications (Faggioli et al., 2023b), e.g., query variant selection (Di Nunzio and Faggioli, 2021; Scells et al., 2018; Thomas et al., 2017), selective query expansion (Amati et al., 2004), IR system configuration selection (Deveaud et al., 2016; Tonellotto et al., 2013), enriching query features for learning-to-rank (Macdonald et al., 2012) , and query-specific pool depth prediction (Ganguly and Yilmaz, 2023) to reduce human relevance judgment costs.
@@
QPP methods can be applied in various domains and scenarios (Meng et al., 2025a; Faggioli et al., 2023b). We are usually concerned with the predicted retrieval quality w.r.t. various IR measures across different scenarios, e.g., our emphasis might be on precision (Faggioli et al., 2023d, 2021a) for conversational search (Lu et al., 2025; Abbasiantaeb et al., 2025) and on recall for legal search (Tomlinson et al., 2007). However, existing QPP approaches typically predict only a single real-valued score that indicates the retrieval quality for a query (Ganguly et al., 2022) and do not require the predicted score to approximate a specific IR evaluation measure (Arabzadeh et al., 2023; Singh et al., 2023; Tao and Wu, 2014; Shtok et al., 2012; Zhou and Croft, 2007).
These properties results in two key limitations: (i) While predicted performance scores have been shown to correlate with some IR evaluation metrics (Datta et al., 2022c; Ganguly et al., 2022), relying on a single value to represent different IR evaluation measures leads to a “one size fits all” approach, which is problematic because the literature shows that some IR metrics do not correlate well and the agreement varies across scenarios and queries (Gupta et al., 2019; Jones et al., 2015). Although some studies train regression-based QPP models to predict a specific IR evaluation measure (Khodabakhsh and Bagheri, 2023; Chen et al., 2022; Datta et al., 2022a; Arabzadeh et al., 2021b; Hashemi et al., 2019), they require training separate models to predict different measures, leading to lots of storage and running costs. (ii) A single-score prediction limits the interpretability of QPP. It is insufficient to explain QPP outputs or to analyze and fix inaccurate QPP results based solely on a single score. We argue that more in-depth and interpretable insights into QPP outputs are required.
@@
We propose a QPP framework using automatically generated relevance judgments (QPP-GenRE), in which we decompose QPP into independent subtasks of automatically predicting the relevance of each item in a ranked list for a given query. QPP-GenRE comes with various advantages: (i) It allows us to directly predict any desired IR evaluation measure at no additional cost, using automatically generated relevance judgments as pseudo-labels. Compared to most existing QPP methods that only outputting a single scalar value as an indicator of a ranker’s overall performance, our method provides a multi-dimensional assessment of system effectiveness by enabling the calculation of various metrics from the same set of relevance judgments. Leveraging predicted relevance judgments, our method allows the computation of metrics such as nDCG@k, Precision@k, reciprocal rank (RR), and so on. This flexibility is particularly advantageous because it allows us to use QPP to predict retrieval quality in terms of a specific evaluation metric that is prioritized in different scenarios. For example, we use predicted relevance judgments to calculate precision-oriented metrics in conversational search, while focusing on recall-oriented metrics for tasks like legal search. (ii) The generated relevance judgments provide an explanation beyond simply gauging how difficult or easy a query is by offering information about why the query is predicted as being difficult or easy; moreover, we can translate the “QPP errors” into easily observable “relevance judgment errors,” e.g., false positives or negatives, informing potential ways of improving QPP quality by fixing observed relevance judgment errors.
Integrating QPP-GenRE with LLM-labeled relevance judgments
@@
QPP-GenRE can be integrated with various approaches for judging relevance. The success of QPP-GenRE depends fundamentally on the accuracy of relevance judgment predictions. Therefore, it is crucial to equip QPP-GenRE with an approach capable of accurately generating relevance judgments. Recently, numerous studies (Takehi et al., 2024; Upadhyay et al., 2024b; Ma et al., 2024; Thomas et al., 2024; Faggioli et al., 2023a; MacAvaney and Soldaini, 2023; Gilardi et al., 2023) have shown the potential effectiveness of using LLMs to generate relevance judgments. Therefore, it is natural to explore equipping QPP-GenRE with LLMs for judging relevance. However, those studies have certain limitations: several authors have prompted commercial LLMs (e.g., ChatGPT, GPT-3.5/4, GPT-4o) to generate relevance judgments (e.g., Chen et al., 2024; Upadhyay et al., 2024c, a; Zhang et al., 2024; Ma et al., 2024; Thomas et al., 2024; Faggioli et al., 2023a); commercial LLMs come with limitations like non-reproducibility, non-deterministic outputs and potential data leakage between pretraining and evaluation data, impeding their value in scientific research (Pradeep et al., 2023b; Zhang et al., 2023c; Pradeep et al., 2023a). Although MacAvaney and Soldaini (2023) prompt small-scale open-source language models (e.g., Flan-T5 (Chung et al., 2022) with 3B parameters) for generating relevance judgments, they focus on a setting wherein the model is already given one relevant item for each query, which does not apply to QPP as we typically do not know any relevant item for a query in advance. In this paper, we focus on the use of open-source LLMs for generating relevance judgments in a realistic setting where we lack prior knowledge of any relevant items for a query. There are only few studies (Zhang et al., 2024; Salemi and Zamani, 2024; Upadhyay et al., 2024a; Khramtsova et al., 2024) attempting to prompt open-source LLMs in this setting.
@@
We face two challenges when using QPP-GenRE for QPP: (i) Predicting IR metrics that not only consider precision but also take recall into account, ideally, entails identifying all relevant items in the entire corpus for a query; however, using an LLM to judge the entire corpus per query is impractical due to the significant computational overhead; (ii) our experiments reveal that directly prompting open-source LLMs in a zero-/few-shot manner yields limited effectiveness in predicting relevance, resulting in limited QPP quality; this aligns with recent findings indicating limited success in prompting open-source LLMs for specific tasks (Qin et al., 2023). Also, incorporating in-context learning examples in few-shot prompting leads to high inference costs (Liu et al., 2022).
@@
Experiments on datasets from the TREC 2019–2022 deep learning (TREC-DL) tracks (Craswell et al., 2022, 2021, 2020, 2019) show that QPP-GenRE achieves state-of-the-art QPP quality in estimating the retrieval quality of a lexical ranker (BM25) and two neural rankers, ANCE (Xiong et al., 2021) and TAS-B (Hofstätter et al., 2021), in terms of RR@10, a precision-oriented IR metric, and nDCG@10, an IR metric considering recall. See Sections 6.1 and 6.2.
@@
Moreover, we conduct an in-depth analysis to investigate the impact of fine-tuning and the choice of LLMs on the quality of generated relevance judgments and QPP. We consider two families of LLMs, Llama and Mistra, with sizes ranging from 1B to 70B, under both few-shot and fine-tuned settings. We find that fine-tuning markedly improves the quality of relevance judgment generation and QPP for all LLMs. In particular, a fine-tuned 3B model (Llama-3.2-3B-Instruct) provides the best trade-off between QPP quality and computational efficiency: it not only significantly outperforms 70B few-shot models, but also achieves QPP quality comparable to that of fine-tuned 7B and 8B models. This suggests that, compared to few-shot prompting, fine-tuning LLMs for relevance prediction can yield higher effectiveness in both relevance prediction and QPP, even with relatively small model sizes; this, in turn, implies that fine-tuning can offer strong performance at lower inference costs. Moreover, the performance of fine-tuned LLMs in terms of judging relevance exceeds that of a commercial LLM (GPT-3.5) (Faggioli et al., 2023a). See Section 7.2.
Additionally, to show QPP-GenRE’s compatibility with other types of relevance prediction methods, we adapt a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a), into a relevance judgment generator by applying a threshold to its re-ranking scores. Our results indicate that QPP-GenRE integrated with RankLLaMA achieves high QPP quality, at the cost of tuning a proper threshold. The high QPP quality achieved by RankLLaMA demonstrates QPP-GenRE’s compatibility with other types of relevance prediction methods. See Section 7.3.
To demonstrate the generalizability of QPP-GenRE to a new domain, we conduct experiments of applying QPP-GenRE to conversational search (Mo et al., 2025; Meng et al., 2025b) in a zero-shot manner. Specifically, we evaluate QPP-GenRE and baselines when predicting the performance of a conversational dense retriever (Yu et al., 2021) on the CAsT-19 (Dalton et al., 2020b) and 20 (Dalton et al., 2020a) datasets. We found that QPP-GenRE consistently outperforms all baselines on both datasets, demonstrating strong generalizability. See Section 7.4.
We also analyze QPP errors based on automatically generated relevance judgments, and provide a case study for a specific example, demonstrating QPP-GenRE’s interpretability. See Section 7.5.
Finally, our computational cost analysis shows that QPP-GenRE shows lower latency than some supervised QPP baselines when predicting multiple measures because multiple measures can be derived from the same set of relevance judgments. Although QPP-GenRE shows higher latency than other QPP baselines when predicting only one metric, QPP-GenRE ’s latency is still 20 times smaller than the state-of-the-art GPT-4-based listwise re-ranker (Sun et al., 2023b). To further enhance the efficiency of QPP-GenRE. We have proposed a relevance judgment caching mechanism. Our experimental results show that the mechanism can reduce LLM calls for relevance prediction by about 30%. Specifically, this mechanism reuses previously predicted relevance judgments for the same query when predicting the performance of new rankers. As a result, this mechanism helps conserve computational resources by avoiding recompute relevance judgments that are shared among multiple rankers. See Section 7.6.
@@
Given QPP-GenRE’s high QPP quality and interpretability, it is well-suited for some knowledge-intensive professional search scenarios, where accurate QPP is prioritized, interpretable QPP results are preferred, and users may have a higher tolerance level for latency than users in web search. Plus, QPP-GenRE can be used to analyze how well a search system performs in offline settings (Faggioli et al., 2023d), where latency is not necessarily an issue.
@@
Another advantage of QPP-GenRE, which makes it applicable to real-world scenarios, is that unlike traditional approaches that depend heavily on the specific properties of rankers, our method is ranker-agnostic. E.g., conventional baselines often rely on score distributions tied to the type of their rankers, making their predictions inherently ranker-dependent. Previous study has demonstrated that the effectiveness of such score-based QPP methods varies across different rankers due to the differences in score distributions produced by each ranker (Faggioli et al., 2023f). In contrast, QPP-GenRE operates on individual query–document pairs and evaluating them independently of a ranker. This eliminates the dependency on specific ranker characteristics and score distributions, ensuring that our framework can be applied generally across various retrieval settings. Furthermore, QPP-GenRE can leverage the reusability of predicted relevance judgments. Since each query–document pair is judged only once, our method allows for predicting the performance of multiple rankers effectively due to their potential overlaps in their top ranked documents. This implies that QPP-GenRE can become more efficient over time as it is used in practice.
@@
To facilitate future research, we release our data, scripts for fine-tuning/inference, sampled demonstration examples for few-shot prompting, and fine-tuned checkpoints of various LLMs at https://github.com/ChuanMeng/QPP-GenRE.
@@
  • We devise an approximation strategy to predict IR measures that account for both precision and recall, avoiding the cost of traversing the entire corpus to identify all relevant items for a query.
  • We fine-tune leading open-source LLMs from the Llama and Mistral families, covering a range of model sizes, for the task of automatically generating relevance judgments. Our results show that fine-tuning much smaller LLMs for relevance judgment prediction can yield more effective relevance prediction and QPP than few-shot prompting with much larger models.
@@ 2.1. Query performance prediction
A large number of unsupervised and supervised post-retrieval QPP methods have been proposed (Carmel and Yom-Tov, 2010) for predicting the performance of lexical rankers, such as query likelihood (Lafferty and Zhai, 2001) and BM25 (Robertson et al., 1995). Unsupervised QPP methods can be classified into clarity-based (Cronen-Townsend et al., 2002), robustness-based (Zhou and Croft, 2007, 2006; Aslam and Pavlu, 2007), coherence-based (Arabzadeh et al., 2021a; Diaz, 2007), and score-based (Tao and Wu, 2014; Shtok et al., 2012; Cummins et al., 2011; Pérez-Iglesias and Araujo, 2010; Zhou and Croft, 2007). More recently, a set of supervised QPP methods have been proposed (Arabzadeh et al., 2021b; Hashemi et al., 2019; Zamani et al., 2018; Datta et al., 2022a, b; Chen et al., 2022; Khodabakhsh and Bagheri, 2023). NeuralQPP (Zamani et al., 2018) and Deep-QPP (Datta et al., 2022a) are optimized from scratch. NQA-QPP (Hashemi et al., 2019) and BERT-QPP (Arabzadeh et al., 2021b) fine-tune BERT (Devlin et al., 2019) to improve QPP effectiveness. Further, Datta et al. (2022c) propose qppBERT-PL, which considers list-wise-document information, while Chen et al. (2022) propose BERT-groupwise-QPP that considers both cross-query and cross-document information. Khodabakhsh and Bagheri (2023) propose a multi-task query performance prediction framework (M-QPPF), learning document ranking and QPP simultaneously.
@@
2.2. Zero/few-shot prompting and parameter-efficient fine-tuning for LLMs
While fine-tuning pre-trained language models has given rise to many state-of-the-art results (Devlin et al., 2019), fully fine-tuning LLMs for a specific task on consumer-level hardware is typically infeasible (Zhu et al., 2023) because of the large number of parameters of LLMs. As a result, there are three prevailing ways to adapt LLMs to a specific task: zero-shot prompting, few-shot prompting (Askari et al., 2025, 2024), a.k.a.  in-context learning (ICL(Dong et al., 2022; Brown et al., 2020), and parameter-efficient fine-tuning (PEFT(Dettmers et al., 2023; Liu et al., 2022; Hu et al., 2021).